Suppose that at least 15.4 eV is needed to free an electron from a particular element, i.e., to ionize the atom. What is the lowest frequency photon which can accomplish this? And, what is the corresponding wavelength for that photon? Give frequency for answer 1 and wavelength for answer 2.
Answer #1

The formula is E = hv
where h = Plancks constant = 6.626*10^-34 J s and v = frequency

E=15.4 ev=15.4*1.6*10^-19=2.464*10^18 j=hv

v=2.464*10^18 /6.63*10^-34=3.71*10^15 Hz

lamda=c/mu=3*10^8/3.71*10^15 =8.07*10^-8 m